OpenStudy (zyberg):

Is there a good way to solve this system of equations: xy + yz = 9; xz + yz = 8; xy + xz = 5?

OpenStudy (zyberg):

I notice that it can be rewritten as: y(x + z) = 9 z(x + y) = 8 x(y + z ) = 5 But I can't find any rational way to solve this.

OpenStudy (sshayer):

add them all 2(xy+yz+xz)=9+8+5=22 xy+yz+xz=22/2=11 now subtract each eq. from this

OpenStudy (3mar):

If I may help?

OpenStudy (sshayer):

i give you hint xy+yz+xz-(xy+yz)=11-9=2 xz=2 similarly yz and xy ten multiply them all xy*yz*xz=? (xyz)^2=? xyz=? now divide one by one e.g., (xyz)/(xy)=? z=? similarly other values.

OpenStudy (loser66):

I just guest!! xy+yz =9 , hence y ( x+z) = 3*3 Assume y = 3, then x + z= 3 for the last one, x(y+z) =5 = 1*5, then assume x =1, then y+z =5, hence z =2 Replace all back, I see (1, 3, 2) satisfies all three equations. :)